OpenStudy (anonymous):

What's the inverse of g(x) = log6 x

OpenStudy (anonymous):

write it with equation please

OpenStudy (anonymous):

This is all the question says Write the inverse of g(x) = log6 x A)G-1(x) = 6x B)G-1(x) =x6 C)G-1(x) = 6y